a cubic foot of water weighs about 62.4 pounds. what must be the diameter of a cylindrical pail 1foot high in order for it to hold the water from 25 pounds f ice?how many square inches of sheet tin are required t make the pan?(neglect waste in cutting and lapping)
Answer by mananth(16949)

a cubic foot of water weighs about 62.4 pounds. what must be the diameter of a cylindrical pail 1foot high in order for it to hold the water from 25 pounds f ice?how many square inches of sheet tin are required t make the pan?(neglect waste in cutting and lapping)
...
62.4 ------------ 1 ft^3
25pounds--------25/62.4 = 0.40 ft^3
...
Volume of cylinder =V = pi*r^2*h
V= pi*r^2*1
0.40/3.14*1= r^2
0.127=r^2
r= 0.35
0.35*12= 4.28 inches
...
Surface area = S = 2*pi*r*h
S=2*pi*4.2*12
S= 322.7 in^2
add area of base= 57.52 in^2
380.22 square inches of sheet.
m.ananth@hotmail.ca